OpenStudy (anonymous):

Help please?? Find the sum of the infinite series 1/3 +4/9 + 16/27+ 64/81+...if it exists

geerky42 (geerky42):

Looks like \[\sum\dfrac{4^n}{3^{n+1}}\]

geerky42 (geerky42):

\[\cdots=\dfrac{1}{3}\sum\dfrac{4^n}{3^n} =\dfrac{1}{3}\sum\left(\dfrac{4}{3}\right)^n\]

geerky42 (geerky42):

Now do you think this series exists?

OpenStudy (anonymous):

yes?

geerky42 (geerky42):

Well, this is geometric series, right?

geerky42 (geerky42):

And we have common ratio greater than 1.

OpenStudy (anonymous):

yeah but I need to know what the sum of the series is

geerky42 (geerky42):

For any finite geometric series, formula is \[\sum_{i=1}^{n-1} r^i = \dfrac{1-r^n}{1-r}\]Right?

OpenStudy (anonymous):

I think so, but what do put in to get the answer? Can you help walk me through the steps? I'm really confused.

geerky42 (geerky42):

Do you know calculus?

OpenStudy (anonymous):

no

geerky42 (geerky42):

Really? ok well just take a look at \[\dfrac{1-r^n}{1-r}\] Since common ratio (r) is greater than 1, what would happen to it if n getting bigger and bigger? \(r^n\) would get bigger and bigger, right?

OpenStudy (anonymous):

yes

geerky42 (geerky42):

well, so when you "get" to infinity. You would have infinity in numerator.

geerky42 (geerky42):

So this series doesn't exist.

geerky42 (geerky42):

Does that make sense?

OpenStudy (anonymous):

Okay thank you <3